(1+6)9÷5 to the second power -4
1
Expert's answer
2019-09-16T10:52:57-0400

[(1+6)×9÷5]2-4 = [7×9÷5]2-4 = [63÷5]2-4 = [12.6]2-4 = 158.76-4


=154.76